# Quantum State Preparation for Probability Distributions with Mirror Symmetry Using Matrix Product States

### v1

@@ 1 Introduction
In various quantum algorithms such as the Monte Carlo method by quantum computer[1, 2, 3], quantum machine learning[4, 5, 6], and simulations for physics[7, 8], quantum state preparation constitutes an important subroutine. The efficiency of quantum state preparation becomes a vital element in determining whether these algorithms can achieve superiority over classical algorithms.
@@
It is known that the preparation of quantum states for arbitrary functions without error requires an exponential depth in quantum circuits[9, 10]. In recent years, several methods have been proposed for state preparation using shallower quantum circuits by tolerating a degree of error. In the context of quantum machine learning, a proposed method involves creating quantum circuits to generate target probability distributions, utilizing quantum Generative Adversarial Networks with parameterized quantum circuits as Generators[11, 12]. Methods utilizing approximations via Fourier transforms have also been proposed[13, 14]. Furthermore, approaches utilizing tensor networks have also been actively researched in recent times[15, 16, 17, 18].
@@
We propose a new state preparation method utilizing the mirror symmetry of probability distributions and tensor networks[16, 17]. We note that functions with less entanglement can be approximated with higher accuracy when using matrix product states. Specifically, in the normal distribution, the portion to the left of the mean monotonically increases and possesses a structure with less entanglement. Therefore, using matrix product states, we could convert the left half of the normal distribution into quantum circuits with higher accuracy. Subsequently, we add a quantum circuit that “duplicates” by the Hadamard gate and inverts this half by CNOT gates. This quantum circuit has enabled us to prepare the entire normal distribution from the quantum circuit that prepares the left half of the normal distribution. Our method achieves a two-order of magnitude improvement in accuracy for the normal distribution compared to methods using existing matrix product states. In addition, we have also confirmed that accuracy improvements can be similarly achieved with distributions that have mirror symmetry other than the normal distribution. Our method features the quantum circuit where most parts consist of gates between nearest-neighbor qubits and exhibit an extremely shallow circuit depth that scales linearly with the number of qubits. If the topology of the quantum computer being used is linear, the fact that CNOT gates act only on nearest-neighbor qubits implies that SWAP gates are not necessary. This makes it a highly useful property for noisy quantum devices. We experimentally showed that when using tensor networks, the accuracy of approximation heavily depends on the bond dimension of the tensor network and is largely independent of the number of qubits. Finally, we demonstrated that we lord the normal distribution into 10 and 20 qubits on a real quantum processor using our designed quantum circuits, achieving high fidelity.
@@ 2 Tensor Networks to Quantum Circuits
Any quantum state of finite dimension can be exactly represented using MPS. If we consider a quantum state as a single tensor and perform singular value decomposition (SVD), we can easily obtain a representation of the quantum state using MPS. However, the bond dimension between the sites of the MPS becomes very large. Generally, a method for exactly converting such MPS into a quantum circuit is unknown.
@@
where $G_{jklm}$ represents the element of the gate $G^{[1]}$, and $j$ and $k$ represents the state of ancilla qubits. Similarly, we can extend $G^{[1]}$ to be a unitary gate. Using these gates $G^{[i]}$, the $U_{{\rm MPD}}$ can be constructed, as shown in Figure 1(a), as follows:
@@
Next, we describe the method when the bond dimension is limited to $\chi>2$. The following approach is based on the idea of reducing the bond dimension by using $U_{{\rm MPD}}$. For instance, we consider the state |$\ket{\Psi^{(4)}}$ when $\chi=4$. We approximate $\chi=2$, and construct $U_{{\rm MPD}}^{(2)}$ using the aforementioned method. Then, using this $U_{{\rm MPD}}^{(2)}$, we disentangle |$\ket{\Psi^{(4)}}$. In other words, we consider $U_{{\rm MPD}}^{(2)}\ket{\Psi^{(4)}}$ to be MPS with approximately $\chi=2$ and approximate $U_{{\rm MPD}}^{(2)}\ket{\Psi^{(4)}}$ again to $\chi=2$ to construct $U_{{\rm MPD}}^{(4)}$. Therefore, it can be considered that $U_{{\rm MPD}}$ of the state |$\ket{\Psi^{(4)}}$ can be approximated as follows:
@@ 3.1 Entanglement and Symmetry
With a lower bond dimension, as shown in Figure 2, it is generally difficult to reproduce any arbitrary probability distribution with high accuracy using MPS. Larger errors occur with lower bond dimensions because the bond dimension is related to the entanglement of the quantum state. This implies that probability distributions with less entanglement will likely have smaller errors than those with larger entanglement.
We consider that by segmenting the probability distribution to reduce entanglement, we could prepare the distribution with greater precision. For example, while the entanglement measure of the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=0.01)$ is $0.237\times 10^{-2}$, the entanglement measure of the left half of the normal distribution is $8.31\times 10^{-6}$. Therefore, it is suggested that the left half of the normal distribution can be approximated with greater precision than the normal distribution. In fact, as shown in Figure 2, it was found that the Kullback-Leibler(KL) divergence is more than two orders of magnitude smaller at lower bond dimensions. Remark that for the entanglement measure, we used the following measure of multiple qubit entanglement proposed in previous research [22]:
@@ 3.2 Result
The results are presented in Figure 4. As shown in Figure 4(a), our approach improved the KL divergence by approximately $10^{-2}$. Additionally, by considering the bond dimension, we found that it is feasible to approximate the probability distribution with the KL divergence of up to the order of $10^{-7}$. Figure 4(b) presents the KL divergence when the bond dimension is constant. It can be observed that the KL divergence is generally independent of the number of qubits and depends primarily on the bond dimension. This suggests that, for approximating the normal distribution, the procedure described in Section 2 can be repeated 11 times to achieve the KL divergence as $\mathcal{O}(10^{-7})$, regardless of the qubit count.
Finally, we demonstrated our methods with a quantum processor ibm_torino using 10 and 20 qubits. We converted learned the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=1)(\mathrm{min}:-4\sqrt{3},\mathrm{max}:4\sqrt{3})$ and measured the final state with 100,000 shots for 10-qubits experiment and 3,000,000 shots for 20-qubits experiment. Several error mitigation methods were utilized by sampler in qiskit runtime primitives. M3 readout error mitigation [25] was performed for the 10-qubit experiment, but not for the 20-qubit one because it timed out. Dynamical decoupling [26, 27] was also enabled.
The results are shown in Fig. 5. Both histograms fit the normal distribution well. Fidelities were 0.879 for 10-qubits and 0.795 for 20-qubits.
Figure 5: Sampled normal distributions by ibm_torino. The number of qubits is 10 qubits for (a) and 20 qubits for (b). The number of samples is 100,000 for one qubits and 3,000,000 for 20 qubits. Fidelity for the probability distributions is 0.879 for 10-qubits and 0.795 for 20-qubits.
@@ 4 Conclusion
We proposed a method that utilizes mirror symmetry to achieve higher precision in preparing the normal distribution when converting from tensor networks to quantum circuits. The depth of this quantum circuit scales linearly with the number of qubits, even when considering the maximum bond dimension. Our method improved the accuracy (KL divergence) by two orders of magnitude compared to existing methods using tensor networks. We showed how bond dimension and the number of qubits influence the precision when preparing the normal distribution using tensor networks. Specifically, we showed that increasing the bond dimension improves accuracy, and the number of qubits has a small impact on accuracy. Furthermore, we confirmed that our method improves accuracy even in distributions with mirror symmetry, aside from the normal distribution. Finally, we implemented our quantum circuit that loads the normal distribution on a real quantum processor. We conducted experiments with 10 qubits and 20 qubits quantum circuits, demonstrating that we could achieve high fidelity in each instance.
@@ 4.1 Comparison with Prior Works
When compared to existing methods using tensor networks[16, 17], the major difference with our approach lies in the accuracy and the CNOT gates between (2) and (3) in Figure 3(a). Regarding accuracy, it is as described above. The CNOT gate acts only between nearest-neighbor qubits in existing tensor network methods. In quantum computers with a linear topology, this characteristic means that SWAP gates are not required. This characteristic, which saves SWAP gates, benefits noisy quantum devices without error correction capabilities. In our method, as shown before point (1) in Figure 3(a), the CNOT gate is applied only between nearest-neighbor qubits. However, the CNOT gates between points (2) and (3) in Figure 3(a) act between the ancilla qubit and all qubits, necessitating SWAP gates. In this respect, our method is at a disadvantage compared to existing methods.
@@
The Grover–Rudolph (GR) state preparation [9] can prepare exact quantum states, but the number of gates and the circuit depths required increases exponentially with the number of qubits. The Kitaev–Webb method [28] allows for similarly accurate state preparation with polynomial circuit depth, but it has been pointed out that for fewer than 15 qubits, it requires a greater circuit depth than the GR method [29]. Additionally, in the Fourier Series Loader (FSL) method [14] using an m-qubit Fourier transform, the GR state preparation is employed to lord the Fourier coefficients, necessitating a circuit depth of $2^{m}$. To prepare quantum states accurately using the FSL method, a moderately large size for m is necessary (for example, around $m=6$, or more). Consequently, this necessitates a significantly large CNOT circuit depth in practice. Moreover, the inverse quantum Fourier transform, which utilizes approximately $\mathcal{O}(n^{2})$ two-qubit gates, also requires a considerably large quantum circuit. A significant disadvantage of these methods is the requirement to use CNOT gates between all qubits, which necessitates many SWAP gates. Since implementing a logical SWAP gate requires three CNOT gates, these methods necessitate a greater number of CNOT gates for implementation. This requirement presents a challenge for execution on noisy quantum devices requiring logical SWAP gates. In contrast, while having less accuracy, our method excels over these methods in terms of shallow CNOT circuit depth and fewer SWAP gate requirements.
@@
Finally, we compare our approach with methods that use machine learning [11, 12, 18]. The advantage of these methods lies in their ability to determine circuit depth and gate count based on the number of parameters. If accurate approximations can be achieved with fewer parameters, these methods can implement shallower quantum circuits compared to our method and other existing methods that depend on the number of qubits. However, in practice, finding a good arrangement of parameters and the parameters themselves is challenging. These studies have been limited to discovering quantum circuits with an accuracy on the order of $\mathcal{O}(10^{-4})$. The fact that the CNOT gates in the quantum circuit act only between nearest-neighbor qubits is beneficial.
@@
In summary, our method offers better precision than approaches using tensor networks or machine learning and maintains a quantum circuit depth comparable to these methods. While it is less precise than methods using GR state preparation, our method significantly benefits from a much shallower quantum circuit. Therefore, our method is particularly powerful for noisy quantum devices with lower CNOT gate fidelity[30, 31].
@@ 4.2 Future Work
We have proposed a method for generating higher-accuracy quantum circuits from matrix product states by utilizing the mirror symmetry of the normal distribution and other functions. In this work, we focused on partially monotonically increasing functions, but extending this approach to other types of functions is a task for the future. Additionally, the precision of our method is contingent upon the accuracy of the conversion from tensor networks to quantum circuits. Hence, proposing methods for more precise conversion is an important challenge. As part of this challenge, it would be intriguing to investigate whether incorporating machine learning [11, 12, 18, 32, 33] into our current method could further enhance accuracy.

### v2

@@ 1 Introduction
In various quantum algorithms such as the Monte Carlo method by quantum computer[1, 2, 3], quantum machine learning[4, 5, 6], and simulations for physics[7, 8], quantum state preparation constitutes a vital subroutine. The efficiency of quantum state preparation becomes vital in determining whether these algorithms can achieve superiority over classical algorithms.
@@
It is known that the preparation of quantum states for arbitrary functions without error requires an exponential depth in quantum circuits[9, 10]. In recent years, several methods have been proposed for state preparation using shallower quantum circuits by tolerating a degree of error. In quantum machine learning, a proposed method involves creating quantum circuits to generate target probability distributions, utilizing quantum Generative Adversarial Networks with parameterized quantum circuits as Generators[11, 12]. Methods utilizing approximations via Fourier transforms have also been proposed[13, 14]. Furthermore, approaches utilizing tensor networks have also been actively researched in recent times[15, 16, 17, 18].
@@
We propose a new state preparation method utilizing the mirror symmetry of probability distributions and tensor networks[16, 17]. We note that functions with less entanglement can be approximated more accurately when using matrix product states. Specifically, in the normal distribution, the portion to the left of the mean monotonically increases and possesses a structure with less entanglement. Therefore, using matrix product states, we could convert the left half of the normal distribution into quantum circuits with higher accuracy. Subsequently, we add a quantum circuit that “duplicates” by the Hadamard gate and inverts this half by CNOT gates. This quantum circuit has enabled us to prepare the entire normal distribution from the quantum circuit that prepares the left half of the normal distribution. Our method achieves a two-order-of-magnitude improvement in normal distribution accuracy compared to methods using existing matrix product states. In addition, we have also confirmed that accuracy improvements can be similarly achieved with distributions with mirror symmetry other than the normal distribution. Our method features a quantum circuit in which most parts consist of gates between nearest-neighbor qubits. The circuit exhibits an extremely shallow circuit depth that scales linearly with the number of qubits. If the topology of the quantum computer being used is linear, the fact that CNOT gates act only on nearest-neighbor qubits implies that SWAP gates are not necessary. This makes it a beneficial property for noisy quantum devices. We experimentally showed that when using tensor networks, the approximation accuracy heavily depends on the bond dimension of the tensor network and is mainly independent of the number of qubits. Finally, we demonstrated that we lord the normal distribution into 10 and 20 qubits on a real quantum processor using our designed quantum circuits, achieving high fidelity.
@@ 2 Tensor Networks to Quantum Circuits
Any quantum state of finite dimension can be exactly represented using MPS. If we consider a quantum state as a single tensor and perform singular value decomposition (SVD), we can quickly obtain a representation of the quantum state using MPS. However, the bond dimension between the sites of the MPS becomes very large. Generally, a method for exactly converting such MPS into a quantum circuit has yet to be discovered.
@@
where $G_{jklm}$ represents the element of the gate $G^{[1]}$, and $j$ and $k$ represents the state of ancilla qubits. Similarly, we can extend $G^{[1]}$ as a unitary gate. Using these gates $G^{[i]}$, the $U_{{\rm MPD}}$ can be constructed, as shown in Figure 1(a), as follows:
@@
Next, we describe the method when the bond dimension is limited to $\chi>2$. The following approach is based on reducing the bond dimension by using $U_{{\rm MPD}}$. For instance, we consider the state |$\ket{\Psi^{(4)}}$ when $\chi=4$. We approximate $\chi=2$, and construct $U_{{\rm MPD}}^{(2)}$ using the aforementioned method. Then, using this $U_{{\rm MPD}}^{(2)}$, we disentangle |$\ket{\Psi^{(4)}}$. In other words, we consider $U_{{\rm MPD}}^{(2)}\ket{\Psi^{(4)}}$ to be MPS with approximately $\chi=2$ and approximate $U_{{\rm MPD}}^{(2)}\ket{\Psi^{(4)}}$ again to $\chi=2$ to construct $U_{{\rm MPD}}^{(4)}$. Therefore, it can be considered that $U_{{\rm MPD}}$ of the state |$\ket{\Psi^{(4)}}$ can be approximated as follows:
@@ 3.1 Entanglement and Symmetry
With a lower bond dimension, as shown in Figure 2, it is generally difficult to reproduce any arbitrary probability distribution with high accuracy using MPS. Lower bond dimensions cause larger errors because the bond dimension is related to the entanglement of the quantum state. This implies that probability distributions with less entanglement will likely have smaller errors than those with larger entanglement.
We consider that segmenting the probability distribution to reduce entanglement could allow us to prepare the distribution with greater precision. For example, while the entanglement measure of the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=0.01)$ is $0.237\times 10^{-2}$, the entanglement measure of the left half of the normal distribution is $8.31\times 10^{-6}$. Therefore, the left half of the normal distribution can be approximated with greater precision than the normal distribution. As shown in Figure 2, it was found that the Kullback-Leibler(KL) divergence is more than two orders of magnitude smaller at lower bond dimensions. Remark that for the entanglement measure, we used the following measure of multiple qubit entanglement proposed in previous research [22]:
@@ 3.2 Result
The results are presented in Figure 4. Figure 4(a) shows that our approach improved the KL divergence by approximately $10^{-2}$. Additionally, by considering the bond dimension, we found it feasible to approximate the probability distribution with the KL divergence of up to the order of $10^{-7}$. Figure 4(b) presents the KL divergence when the bond dimension is constant. It can be observed that the KL divergence is generally independent of the number of qubits and depends primarily on the bond dimension. This suggests that, for approximating the normal distribution, the procedure described in Section 2 can be repeated 11 times to achieve the KL divergence as $\mathcal{O}(10^{-7})$, regardless of the qubit count.
Finally, we demonstrated our methods with a quantum processor ibm_torino using 10 and 20 qubits. We converted and learned the normal distribution $\mathcal{N}(\mu=0,\sigma^{2}=1)(\mathrm{min}:-4\sqrt{3},\mathrm{max}:4\sqrt{3})$ and measured the final state with 100,000 shots for 10-qubits experiment and 3,000,000 shots for 20-qubits experiment. Several error mitigation methods were utilized by the sampler in qiskit runtime primitives. M3 readout error mitigation [25] was performed for the 10-qubit experiment but not for the 20-qubit one because it timed out. Dynamical decoupling [26, 27] was also enabled.
The results are shown in Fig. 5. Both histograms fit the normal distribution well. Fidelities were 0.879 for 10 qubits and 0.795 for 20 qubits.
Figure 5: Sampled normal distributions by ibm_torino. The qubits are 10 qubits for (a) and 20 qubits for (b). The number of samples is 100,000 for 10 qubits and 3,000,000 for 20 qubits. The fidelity for the probability distributions is 0.879 for 10 qubits and 0.795 for 20 qubits.
@@ 4 Conclusion
We proposed a method that utilizes mirror symmetry to achieve higher precision in preparing the normal distribution when converting from tensor networks to quantum circuits. The depth of this quantum circuit scales linearly with the number of qubits, even when considering the maximum bond dimension. Our method improved the accuracy (KL divergence) by two orders of magnitude compared to existing methods using tensor networks. We showed how bond dimension and the number of qubits influence the precision when preparing the normal distribution using tensor networks. Specifically, we showed that increasing the bond dimension improves accuracy, and the number of qubits has a negligible impact on accuracy. Furthermore, we confirmed that our method improves accuracy even in distributions with mirror symmetry, in addition to the normal distribution. Finally, we implemented our quantum circuit that loads the normal distribution on a real quantum processor. We conducted experiments with 10 qubits and 20 qubits quantum circuits, demonstrating that we could achieve high fidelity in each instance.
@@ 4.1 Comparison with Prior Works
When compared to existing methods using tensor networks[16, 17], the significant difference with our approach lies in the accuracy and the CNOT gates between (2) and (3) in Figure 3(a). Regarding accuracy, it is as described above. The CNOT gate acts only between nearest-neighbor qubits in existing tensor network methods. This characteristic in quantum computers with a linear topology means that SWAP gates are not required. This characteristic, which saves SWAP gates, benefits noisy quantum devices without error correction capabilities. In our method, as shown before point (1) in Figure 3(a), the CNOT gate is applied only between nearest-neighbor qubits. However, the CNOT gates between points (2) and (3) in Figure 3(a) act between the ancilla qubit and all qubits, necessitating SWAP gates. In this respect, our method is at a disadvantage compared to existing methods.
@@
The Grover–Rudolph (GR) state preparation [9] can prepare exact quantum states, but the number of gates and the circuit depths required increases exponentially with the number of qubits. The Kitaev–Webb method [28] allows for similarly accurate state preparation with polynomial circuit depth. However, it has been pointed out that for fewer than 15 qubits, it requires a greater circuit depth than the GR method [29]. Additionally, in the Fourier Series Loader (FSL) method [14] using an m-qubit Fourier transform, the GR state preparation is employed to lord the Fourier coefficients, necessitating a circuit depth of $2^{m}$. To prepare quantum states accurately using the FSL method, a moderately large size for m is necessary (for example, around $m=6$, or more). Consequently, this necessitates a significantly large CNOT circuit depth in practice. Moreover, the inverse quantum Fourier transform, which utilizes approximately $\mathcal{O}(n^{2})$ two-qubit gates, also requires a considerably sizeable quantum circuit. A significant disadvantage of these methods is the requirement to use CNOT gates between all qubits, which necessitates many SWAP gates. Since implementing a logical SWAP gate requires three CNOT gates, these methods necessitate a more significant number of CNOT gates for implementation. This requirement presents a challenge for executing noisy quantum devices requiring logical SWAP gates. In contrast, while our method is less accurate, it excels over these methods in terms of shallow CNOT circuit depth and fewer SWAP gate requirements.
@@
Finally, we compare our approach with methods that use machine learning [11, 12, 18]. These methods’ advantage is their ability to determine circuit depth and gate count based on the number of parameters. If accurate approximations can be achieved with fewer parameters, these methods can implement shallower quantum circuits compared to our method and other existing methods that depend on the number of qubits. However, finding a suitable arrangement of parameters and the parameters themselves is challenging in practice. These studies have been limited to discovering quantum circuits with an accuracy on the order of $\mathcal{O}(10^{-4})$. The fact that the CNOT gates in the quantum circuit act only between nearest-neighbor qubits is beneficial.
@@
In summary, our method offers better precision than approaches using tensor networks or machine learning and maintains a quantum circuit depth comparable to these methods. While it is less precise than GR state preparation methods, our method significantly benefits from a much shallower quantum circuit. Therefore, our method is compelling for noisy quantum devices with lower CNOT gate fidelity[30, 31].
@@ 4.2 Future Work
We have proposed a method for generating higher-accuracy quantum circuits from matrix product states by utilizing the mirror symmetry of the normal distribution and other functions. In this work, we focused on partially monotonically increasing functions, but extending this approach to other types of functions is a task for the future. Additionally, the precision of our method is contingent upon the accuracy of the conversion from tensor networks to quantum circuits. Hence, proposing methods for more precise conversion is a significant challenge. As part of this challenge, it would be intriguing to investigate whether incorporating machine learning [11, 12, 18, 32, 33] into our current method could further enhance accuracy.